OpenStudy (gabby1214):

solve for x -5/2x <= -10 the solution is x >= ?

OpenStudy (anonymous):

is it \[x-\frac{ 5 }{ 2 }x \le -10\] ?

OpenStudy (gabby1214):

yes

OpenStudy (anonymous):

\[2x-5x \le -20\] \[-3x \le -20,x \ge \frac{ -20 }{ -3 },or ~x \ge \frac{ 20 }{ 3 }\]

OpenStudy (gabby1214):

thank you!

OpenStudy (anonymous):

yw